OpenStudy (anonymous):

how to convert 18 years to seconds

OpenStudy (anonymous):

There are 365 days in a year. 24 hours in a day. 60 minutes in an hour. And 60 seconds in a minute. With this information do you think you could solve this?

OpenStudy (anonymous):

18*365*24*60*60

OpenStudy (anonymous):

yes thanks! i get it

OpenStudy (anonymous):

18*365=6570 24*6570=157680 157680*60=9460800 9460800*60=567648000 Do you see what me and @surjithayer did?

OpenStudy (anonymous):

yes it helped

OpenStudy (anonymous):

I'm glad we could assist!